The mean of 5 observations is 15 and the


variance is 9. If two more observations


having values – 3 and 10 are combined


with these 5 observations, what will be the


variance of the


7 observations ?

1
Expert's answer
2022-01-11T15:16:11-0500

for 5 observations:

"\sum x_i=\mu n=15\cdot5=75"

for 7 observations:

"\sum x_i=75-3+10=82"

then, the new mean is

"\mu_1=82/7"


variance for 5 observations:

"\sigma^2=\sum x_i^2/5-\mu^2=\sum x_i^2/5-15^2=9"

then

"\sum x_i^2=5\cdot(9+15^2)=1170"


for 7 observations:

"\sum x_i^2=1170+3^2+10^2=1279"

new variance:

"\sigma_1^2=1279/7-\mu_1^2=1279/7-82^2/7^2=45.49"